# Beyond LLMs: An Exploration of Small Open-source Language Models in Logging Statement Generation

### v1

@@ 3.2. Studied Models
  • LLaMA (Grattafiori et al., 2024) is Meta’s latest LLM and refines the Llama 2 framework. It stands as a prominent open source LLM used in numerous software applications. Trained on an extensive and varied dataset far surpassing its predecessor, Llama 3 exhibits significantly improved proficiency in reasoning, code generation, and instruction adherence.
@@
  • Qwen2.5-coder (Hui et al., 2024) is a code-specialized version of the Qwen2 (Yang et al., 2024a) family, which inherits Qwen’s multilingual capabilities and architectural improvement. While demonstrating strong and comprehensive coding abilities, it also possesses good general and mathematical skills.
@@ 3.3. Baselines
To evaluate SOLMs performance, we selected the baselines by conducting a literature review of relevant papers published in SE venues. From this, we find the following four methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2024), UniLog (Xu et al., 2024), and FastLog (Xie et al., 2024). Additionally, we examine several notable LLM baselines, including general-purpose LLMs (Claude3.7-sonnet, GPT4o, LLAMA-405b), and a code-specific LLM (Deepseek-coder-v3). In relation to the methodologies for prompting, we derive our approach from the study (Gao et al., 2024) and incorporate four different strategies: instruction prompting (base), in-context learning (ICL), retrieval-augmented generation (RAG) and chain of thought (CoT). The instruction prompting strategy involves directly prompting LLMs to generate logging statements using identical inputs as those provided to SOLMs, without any supplementary data. The ICL approach consists of providing one random example before the main query to assist the model in grasping the nature of the task more effectively. In the RAG approach, rather than selecting example randomly, we retrieve precise examples from the valid dataset to accommodate various input samples. We specifically employ the BM25 algorithm to identify the case within the valid dataset that most closely resembles the query sample in terms of similarity. The details of the prompts are shown in Figure 3.
@@ 3.4. Strategies for Parameter-Efficient Fine-Tuning
We examined how different PEFT strategies influence the performance of SOLMs when generating automated logging statements.
@@
3.5. Evaluation Metrics
@@ 3.5.1. Logging point
Position Accuracy (PA): To quantify PA, we compare the predicted locations of logging statements against their ground truth positions in the source code. This metric is formally defined as the ratio of correctly positioned logging statements ($LS_{position\_correct}$) to the total number of logging statements ($LS_{all}$), expressed as $\frac{LS_{position\_correct}}{LS_{all}}$.
3.5.2. Logging level
@@
We use the Level Accuracy (LA) and Average Ordinal Distance Score (AOD) for evaluating the prediction of logging levels. Given the significant semantic differences between these levels and their implications for system monitoring and maintenance, we rigorously assess LA by comparing predicted log levels against their ground truth values in the source code. This metric is formally defined as the ratio of correctly predicted log levels ($LS_{level\_correct}$) to the total number of logging statements ($LS_{all}$), expressed as $\frac{LS_{level\_correct}}{LS_{all}}$. AOD evaluates how closely the current logging level aligns with the recommended logging level for each specific logging statement, as detailed in (Li et al., 2021a). The formula to calculate AOD is given by: $AOD=\frac{\sum_{i=1}^{N}(1-\frac{Dis(a_{i},s_{i})}{MaxDis(a_{i})})}{N}$, where $N$ represents the total number of logging statements in consideration. The term $MaxDis(a_{i})$ is used to denote the maximum potential distance for the actual log level $a_{i}$.
3.5.3. Static logging text
@@
Same as previous study (Li et al., 2024a; Mastropaolo et al., 2022; Ding et al., 2022) , our evaluation of static logging texts is conducted through the application of two metrics commonly employed in the domain of machine translation: BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004). These metrics, grounded in n-gram analysis, are instrumental in assessing the degree of similarity between log messages that are generated computationally and those authored by developers. They provide a normalized score continuum from 0 to 1, with elevated scores indicative of a closer resemblance. In our methodology, we specifically implement various forms of these metrics, identified as BLEU-4 and ROUGE-L.
3.5.4. Dynamic logging variables
@@
We use Precisely Match Rate (PMR) and F1 to evaluate dynamic logging variables. PMR ensures consistency in the capture of variable runtime data—a critical aspect of log effectiveness. We extract dynamic variables from both reference implementations and predicted logging statements, then perform exact matching to evaluate correspondence. PMR is formally defined as the ratio of exactly matched dynamic variables ($LS_{variable\_correct}$) to the total number of logging statements ($LS_{all}$), expressed as $\frac{LS_{variable\_correct}}{LS_{all}}$. Moreover, consider each logging statement and let us define $S_{ud}$ as the set of variables included in the generated logging statement, while $S_{gt}$ pertains to the set of variables present in the actual logging statement. In our analysis, we determine and present the following metrics: the precision, which is the ratio of variables from the updates that accurately match those in the actual logging ($precision=\frac{S_{ud}\cap S_{gt}}{S_{ud}}$); the recall, which indicates the fraction of actual variables correctly anticipated by the model ($recall=\frac{S_{ud}\cap S_{gt}}{S_{gt}}$); and, lastly, their harmonic mean, expressed as the F1 score ($F_{1}=2*\frac{Precision*Recall}{Precision+Recall}$(Li et al., 2024a).
3.6. LLM-as-a-judger
@@
Specifically, we select three LLMs recognized for their superior performance in code-related tasks: Claude3.7-Sonnet, Deepseek-coder-v3, and GPT-4o. We choose this model due to their robust code comprehension and generation capabilities, which are critical for assessing the nuanced quality of logging statements. The evaluation process involves providing each LLM judge with the input code context, the generated logging statement from model, and the corresponding ground truth logging statement. The judges assign scores ranging from 0 to 3, where higher scores indicate greater alignment with the ground truth in terms of logging point accuracy, level appropriateness, static text quality, and dynamic variable correctness. To ensure consistency and reliability, we develop a comprehensive scoring guideline, which outline specific criteria for evaluating each component of the logging statement. These criteria address syntactic accuracy, semantic relevance, and contextual appropriateness, mitigating the limitations of traditional metrics like BLEU-4 and ROUGE-L.
@@ 3.7. Implimentation details
4. Experimantal Results
4.1. RQ1: What are the most effective interaction strategies for using SOLMs in logging generation?
@@
Approach. In addressing this question, our study is focused on evaluating the effectiveness of current prompting techniques within the domain of log generation. We specifically analyze, based on the definitions laid out in Section 3.3, the effectiveness of several prompting strategies: basic instruction prompting (base), in-context learning (ICL), retrieval-augmented generation (RAG), and chain-of-thought (CoT). The specific details regarding the prompt templates used can be found in Figure 3. To strengthen the generalizability of our findings, we employ multiple 7B instruction-following models, namely LLaMA, Mistral, CodeLlama, and Qwen2.5-Coder, all in their original, un-fine-tuned configurations.
Figure 3. The prompt template for automated logging statement generation.
@@
Results. The quantitative evaluation comparing the four prompting techniques across the four selected 7B SOLMs is presented in Table 2. Our analysis of these results yields two primary findings concerning the performance of un-fine-tuned models and the efficacy of different interaction strategies for automated logging generation.
@@
First, we observe a distinct performance disparity between the general-purpose instruction-following models (LLaMA, Mistral) and the code-specific models (CodeLlama, Qwen2.5Coder) when used without any task-specific fine-tuning. The general-purpose models generally demonstrate superior performance on this task. For instance, LLaMA and Mistral achieve peak PA scores of 21.01 and 18.15 respectively (both using RAG), substantially higher than the peak PA scores achieved by CodeLlama (11.62 with ICL) and Qwen2.5Coder (13.25 with CoT). We attribute this overall trend to the relatively weaker instruction-following capabilities observed in the un-fine-tuned code-specific models for this specific task. We noted a higher tendency for CodeLlama and Qwen2.5Coder to refuse the prompt or return non-parsable empty responses compared to LLaMA and Mistral, which negatively impacts their effectiveness in this experimental setting. This suggests that, without fine-tuning, the broader instruction comprehension of general-purpose models may be more advantageous for complex code-related generation tasks like automated logging statement generation than the specialized but potentially more rigid capabilities of un-fine-tuned code models.
@@
Findings 1: Without fine-tuning, general-purpose models outperform code-specific models for automated logging statement generation due to the former’s better instruction following ability.
@@
Second, RAG emerges as the most effective prompting technique for enhancing automated logging generation genertaion performance. While other techniques occasionally yield the top score for an isolated metric, RAG demonstrates the most significant and robust improvements across the majority of metrics and models. Notably, for Mistral, RAG achieves the highest scores across all evaluated metrics, including PA (18.15), F1 (40.21), and ROUGE-L (34.74). For LLaMA, RAG secures the top performance in PA (21.01), PMR (41.68), F1 (46.28), BLEU-4 (15.95), and ROUGE-L (36.73). For CodeLlama and Qwen2.5Coder, RAG generally leads to substantial gains over the baseline, particularly in metrics like F1 (CodeLlama: 39.01, Qwen2.5Coder: 49.28) and ROUGE-L (CodeLlama: 34.36, Qwen2.5Coder: 37.57). These results strongly indicate that providing relevant contextual information retrieved from a knowledge base significantly aids the SOLMs in accurately determining where to log, what variables to include, and formulating appropriate log messages, making RAG a highly promising strategy.
@@
Qwen2.5Coder base 4.30 58.91 83.14 27.91 29.55 8.61 27.13
@@
4.2. RQ2: What’s the best strategy using SOLMs for automated logging statement generation?
@@
(RQ2.1) Which PEFT technique yields the optimal performance for automoated logging statement generation using SOLMs? While SOLMs are more compact than their larger counterparts, fully fine-tuning them for specific downstream tasks like automated logging statement generation can still be computationally demanding and may risk overfitting, especially with limited task-specific data. PEFT methodologies have emerged as a compelling solution, enabling adaptation by updating only a small fraction of the model’s parameters or by adding a small set of new, trainable parameters. This significantly reduces computational costs and the risk of catastrophic forgetting of the model’s pre-trained knowledge. However, a diverse range of PEFT techniques exists, each employing different mechanisms to inject task-specific information into the SOLM. The relative efficacy of these techniques can vary depending on the nature of the downstream task. For logging statement generation, which involves understanding code context, identifying appropriate logging locations, and generating relevant log messages, it is unclear which PEFT strategy offers the optimal balance of performance and efficiency. Therefore, we conduct an evaluation of the performance of SOLMs fine-tuned with various prominent PEFT techniques to ascertain which techniques yield superior performance outcomes.
(RQ2.2) How does the size of SOLMs impact the performance-resource trade-offs in automated logging statement generation? The number of parameters in a language model is a critical factor that typically dictates its representational ability, learning capacity, and consequently, its performance on downstream tasks. Although we focus on SOLMs, there is still considerable variation in size within this class. Larger models might capture more complex code patterns and nuances relevant to logging, potentially leading to higher accuracy, but they also incur greater computational costs during fine-tuning and inference. Therefore, evaluating the impact of model size is essential to understand the performance-resource trade-offs specific to automated logging statement generation.
(RQ2.3) Does the instruct variant of a SOLM outperform its base counterpart for automated logging statement generation? The distinction between using a ‘base’ pre-trained model versus an ‘instruct’ version of an SOLM presents another critical strategic choice. Instruct models are specifically optimized to understand and respond to user prompts and instructions, which could be beneficial for a directed task. They might more readily produce outputs that adhere to desired formats or incorporate specified information. However, this instruction tuning is often general-purpose. Base models, on the other hand, represent the raw capabilities learned during pre-training and might offer greater plasticity when fine-tuned on a highly specific downstream task like ours. It is plausible that fine-tuning a base model directly on logging data could lead to a more specialized and potentially more effective model, whereas an instruct model might carry over behaviors from its instruction-tuning phase that are not optimally aligned with logging generation. Therefore, we aim to clarify which model mode serves as a better foundation for this task.
@@
For RQ2.2, focusing on the influence of model size, we selected the Qwen2.5 Coder model series. This choice is driven by the public availability of multiple versions within the same model family, specifically those with 0.5B, 1.5B, 3B, 7B, and 14B parameters, enabling a controlled comparison. Based on the findings from RQ2.1 where LoRA demonstrated superior performance among the PEFT techniques, we exclusively employ LoRA for fine-tuning across these different model sizes. Furthermore, we evaluat each size using both the basic prompt (‘base’) and the RAG-enhanced prompt (‘RAG’), in order to explore whether the effectiveness of RAG varies with model scale, particularly to assess the RAG capabilities of the smaller SOLMs.
To address RQ2.3, we compare the performance of ‘base’ models against their ‘instruct’ counterparts for automated logging statement generation. We select the Mistral 7B model, available in both base and instruct variants, for a controlled comparison, as it got the best perfomance in RQ2.1. Both model modes are fine-tuned using LoRA and employed the RAG-enhanced prompt, consistent with the result of RQ2.1 and RQ2.2. The fine-tuning process spanned five epochs, and performance are evaluated using metrics including Reject Rate, PA, LA, AOD, PMR, F1, BLEU-4, and ROUGE-L. The Reject Rate, defined as the percentage of test set examples rejected by the fine-tuned model during inference due to misalignment with learned task-specific criteria, quantifies the model’s selectivity, reflecting its ability to adhere to prompt instructions and generate valid logging statements.
Table 3. Performance Comparison of PEFT Techniques for SOLMs in Automated Logging Statement Generation.
@@
Results. (RQ2.1) Table 3 shows the performance comparison of PEFT techiniques for SOLMs in automated logging statement generation. We find that all evaluated PEFT methods consistently improve performance over the baseline across all SOLMs and most metrics. For instance, looking at the prediction accuracy, QLoRA fine-tuning increased PA from 13.32 to 45.27 for LLaMA, from 12.85 to 61.90 for Mistral, from 6.99 to 58.97 for CodeLlama, and from 4.30 to 60.21 for Qwen2.5Coder. Similar substantial gains are observed across other metrics like F1 score for variable prediction and BLEU-4/ROUGE-L for statement generation, indicating that fine-tuning with parameter-efficient techniques, is crucial for adapting SOLMs to this specific task.
Furthermore, LoRA generally emerges as the most effective PEFT methodology, closely followed by QLoRA. LoRA achieves the highest scores for the majority of metrics across all four models. For example, with LoRA, LLaMA gets the top PA (56.84), PMR (50.15), F1 (58.22), BLEU-4 (19.84), and ROUGE-L (40.89). Similarly, LoRA leads to the best PA (63.97), LA (69.50), BLEU-4 (23.40), and ROUGE-L (45.73) for Mistral. CodeLlama shows LoRA as the top performer across all metrics. For Qwen2.5Coder, LoRA is also dominant, securing the best PA (62.40), LA (68.46), AOD (88.82), PMR (52.24), BLEU-4 (23.04), and ROUGE-L (44.96). QLoRA consistently performs very competitively, often achieving the second-best results or even surpassing LoRA in a few specific instances (e.g., F1 for Qwen2.5Coder, PMR and F1 for Mistral). Prompt Tuning shows some strength, particularly for ‘level’ prediction with LLaMA, while Prefix Tuning, though an improvement over the baseline, is generally outperformed by LoRA, QLoRA, and Prompt Tuning.
@@
Findings 3: Fine-tuning with PEFT techniques significantly enhances SOLMs performance for automated logging statement generation, with LoRA demonstrating the most consistent and superior results across the evaluated models and metrics.
@@
(b) Qwen2.5 Coder of Different Sizes
@@
Table 4. Impact of Model Size on Performance and Resource Usage for Automated Logging Statement Generation.
@@
(RQ2.2) Table 4 shows the impact of model size on performance and resource usage for automated logging statement generation. We can find that increasing model size generally leads to improved performance in automated logging statement generation, particularly for models 3B and larger. Across almost all metrics, there is a discernible improvement as the model parameter count increases from 0.5B to 14B. For example, prediction accuracy (PA) improves from 21.38 (0.5B) to 66.20 (14B), and ROUGE-L scores for text generation increase from 39.55 (0.5B) to 47.22 (14B). The 14B model consistently outperforms all smaller variants, and the 7B model also shows strong performance.

### v2

@@ 3.2. Studied Models
  • LLaMA3 (Grattafiori et al., 2024) is Meta’s latest LLM and refines the LLaMA 2 framework. It stands as a prominent open source LLM used in numerous software applications. Trained on an extensive and varied dataset far surpassing its predecessor, LLaMA 3 exhibits significantly improved proficiency in reasoning, code generation, and instruction adherence.
@@
  • Qwen2.5-coder (Hui et al., 2024) is a code-specialized version of the Qwen2.5 (Yang et al., 2024a) family, which inherits Qwen’s multilingual capabilities and architectural improvement. While demonstrating strong and comprehensive coding abilities, it also possesses good general and mathematical skills.
@@ 3.3. Baselines
To evaluate SOLMs performance, we select the baselines by conducting a literature review of relevant papers published in SE venues. From this, we find the following methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2024), UniLog (Xu et al., 2024), and FastLog (Xie et al., 2024). Additionally, we examine several LLM baselines, including general-purpose LLMs (Claude3.7-sonnet, GPT4o, LLaMA3.1-405b), and a code-specific LLM (Deepseek-coder-v3). In relation to the methodologies for prompting, we derive our approach from the study (Gao et al., 2024) and incorporate four different strategies: instruction prompting (base), in-context learning (ICL), retrieval-augmented generation (RAG) and chain of thought (CoT). The instruction prompting strategy involves directly prompting LLMs to generate logging statements using identical inputs as those provided to SOLMs, without any supplementary data. The ICL approach consists of providing one random example before the main query to assist the model in grasping the nature of the task more effectively. In the RAG approach, rather than selecting example randomly, we retrieve precise examples from the valid dataset to accommodate various input samples. We specifically employ the BM25 algorithm to identify the case within the valid dataset that most closely resembles the query sample in terms of similarity. The details of the prompts are shown in Figure 3.
@@ 3.4. Strategies for Parameter-Efficient Fine-Tuning
We examine how the following PEFT strategies influence the performance of SOLMs when automated logging.
@@
3.5. Evaluation method
3.5.1. Tradictional evaluating metrics
@@
Logging point: We use position accuracy (PA) to evaluate the performance of logging location. To quantify PA, we compare the predicted locations of logging statements against their ground truth positions in the source code. This metric is formally defined as the ratio of correctly positioned logging statements ($LS_{position\_correct}$) to the total number of logging statements ($LS_{all}$), expressed as $\frac{LS_{position\_correct}}{LS_{all}}$.
@@
Logging level: We use the Level Accuracy (LA) and Average Ordinal Distance Score (AOD) for evaluating the prediction of logging levels. Given the significant semantic differences between these levels and their implications for system monitoring and maintenance, we rigorously assess LA by comparing predicted log levels against their ground truth values in the source code. This metric is formally defined as the ratio of correctly predicted log levels ($LS_{level\_correct}$) to the total number of logging statements ($LS_{all}$), expressed as $\frac{LS_{level\_correct}}{LS_{all}}$. AOD evaluates how closely the current logging level aligns with the recommended logging level for each specific logging statement, as detailed in (Li et al., 2021a). The formula to calculate AOD is given by: $AOD=\frac{\sum_{i=1}^{N}(1-\frac{Dis(a_{i},s_{i})}{MaxDis(a_{i})})}{N}$, where $N$ represents the total number of logging statements in consideration. The term $MaxDis(a_{i})$ is used to denote the maximum potential distance for the actual log level $a_{i}$.
@@
Static logging text: Same as previous study (Li et al., 2024a; Mastropaolo et al., 2022; Ding et al., 2022) , our evaluation of static logging texts is conducted through the application of two metrics commonly employed in the domain of machine translation: BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004). These metrics, grounded in n-gram analysis, are instrumental in assessing the degree of similarity between log messages that are generated computationally and those authored by developers. They provide a normalized score continuum from 0 to 1, with elevated scores indicative of a closer resemblance. In our methodology, we specifically implement various forms of these metrics, identified as BLEU-4 and ROUGE-L.
@@
Dynamic logging variables: We use Precisely Match Rate (PMR) and F1 to evaluate dynamic logging variables. PMR ensures consistency in the capture of variable runtime data—a critical aspect of log effectiveness. We extract dynamic variables from both reference implementations and predicted logging statements, then perform exact matching to evaluate correspondence. PMR is formally defined as the ratio of exactly matched dynamic variables ($LS_{variable\_correct}$) to the total number of logging statements ($LS_{all}$), expressed as $\frac{LS_{variable\_correct}}{LS_{all}}$. Moreover, consider each logging statement and let us define $S_{ud}$ as the set of variables included in the generated logging statement, while $S_{gt}$ pertains to the set of variables present in the actual logging statement. In our analysis, we determine and present the following metrics: the precision, which is the ratio of variables from the updates that accurately match those in the actual logging ($precision=\frac{S_{ud}\cap S_{gt}}{S_{ud}}$); the recall, which indicates the fraction of actual variables correctly anticipated by the model ($recall=\frac{S_{ud}\cap S_{gt}}{S_{gt}}$); and, lastly, their harmonic mean, expressed as the F1 score ($F_{1}=2*\frac{Precision*Recall}{Precision+Recall}$(Li et al., 2024a).
3.5.2. LLM-as-a-judge
@@
Specifically, we select three LLMs recognized for their superior performance in code-related tasks: Claude3.7-Sonnet, Deepseek-coder-v3, and GPT-4o. We choose this model due to their robust code comprehension and generation capabilities, which are critical for assessing the nuanced quality of logging statements. The evaluation process involves providing each LLM judge with the input code context, the generated logging statement from model, and the corresponding ground truth logging statement. The judges assign scores ranging from 0 to 3, where higher scores indicate greater alignment with the ground truth in terms of logging point accuracy, level appropriateness, static text quality, and dynamic variable correctness. For each generated logging statement, the final score is the average of the scores provided by these three LLM judges. To ensure consistency and reliability, we develop a comprehensive scoring guideline, which outline specific criteria for evaluating each component of the logging statement. These criteria address syntactic accuracy, semantic relevance, and contextual appropriateness, mitigating the limitations of traditional metrics like BLEU-4 and ROUGE-L.
@@ 3.6. Implimentation Details
4. Empircial Study Results
4.1. RQ1: What are the most effective prompting strategies for using SOLMs in automated logging?
@@
Approach. In addressing this question, our study is focused on evaluating the effectiveness of current prompting techniques within the domain of log generation. We specifically analyze, based on the definitions laid out in Section 3.3, the effectiveness of several prompting strategies: basic instruction prompting (base), in-context learning (ICL), retrieval-augmented generation (RAG), and chain-of-thought (CoT). The specific details regarding the prompt templates used can be found in Figure 3. To strengthen the generalizability of our findings, we employ multiple 7B instruction-following models, namely LLaMA3, Mistral, CodeLlama, and Qwen2.5-Coder, all in their original configurations.
Figure 3. The prompt template for automated logging.
@@
Results. The quantitative evaluation comparing the four prompting techniques across the four selected 7B SOLMs is presented in Table 2. Our analysis of these results yields two primary findings concerning the performance of un-fine-tuned models and the efficacy of different prompt strategies for automated logging generation.
@@
First, we observe a distinct performance disparity between the general-purpose instruction-following models (LLaMA3, Mistral) and the code-specific models (CodeLlama, Qwen2.5-coder) when used without any task-specific fine-tuning, i.e., the general-purpose models generally demonstrate superior performance on logging. For instance, LLaMA3 and Mistral achieve peak PA scores of 21.01 and 18.15, respectively (both using RAG), substantially higher than the peak PA scores achieved by CodeLlama (11.62 with ICL) and Qwen2.5-coder (13.25 with CoT). We attribute this trend to the weaker instruction-following capabilities in the original code-specific models. Furthermore, upon analyzing the failure cases, we noted a higher tendency for CodeLlama and Qwen2.5-coder to refuse the prompt or return empty responses compared to LLaMA3 and Mistral, which negatively impacts their effectiveness in this experimental setting. This suggests that, without fine-tuning, the broader instruction comprehension of general-purpose models may be more advantageous for automated logging tasks.
@@
Findings 1: Without fine-tuning, general-purpose models outperform code-specific models for automated logging due to their better instruction-following ability.
@@
Second, RAG emerges as the most effective prompting technique for enhancing automated logging generation genertaion performance. While other techniques occasionally yield the top score for an isolated metric, RAG demonstrates the most significant and robust improvements across the majority of metrics and models. Notably, for Mistral, RAG achieves the highest scores across all evaluated metrics, including PA (18.15), F1 (40.21), and ROUGE-L (34.74). For LLaMA3, RAG secures the top performance in PA (21.01), PMR (41.68), F1 (46.28), BLEU-4 (15.95), and ROUGE-L (36.73). For CodeLlama and Qwen2.5-coder, RAG generally leads to substantial gains over the baseline, particularly in metrics like F1 (CodeLlama: 39.01, Qwen2.5-coder: 49.28) and ROUGE-L (CodeLlama: 34.36, Qwen2.5-coder: 37.57). These results strongly indicate that providing relevant contextual information retrieved from a knowledge base significantly aids the SOLMs in accurately determining where to log, what variables to include, and formulating appropriate log messages, making RAG a highly promising strategy.
@@
Qwen2.5-coder base 4.30 58.91 83.14 27.91 29.55 8.61 27.13
@@
4.2. RQ2: What’s the best tuning strategy using SOLMs for automated logging?
@@
(RQ2.1) Which PEFT technique yields the optimal performance for automoated logging statement generation using SOLMs? While SOLMs are more compact than larger LLMs, fully fine-tuning them for specific downstream tasks like automated logging can still be computationally demanding and may risk overfitting on task-specific data. PEFT methodologies have emerged as a compelling solution, enabling adaptation by updating only a small fraction of the model’s parameters or by adding a small set of new, trainable parameters. This significantly reduces computational costs and avoids catastrophic forgetting of the model’s pre-trained knowledge. However, a diverse range of PEFT techniques exists, each employing different mechanisms to inject task-specific information into the SOLM. For automated logging, which involves understanding code context, identifying appropriate logging locations, and generating relevant log messages, it is unclear which PEFT strategy offers the optimal balance of performance and efficiency. Therefore, we conduct an evaluation of the performance of SOLMs fine-tuned with various prominent PEFT techniques to ascertain which techniques yield superior performance outcomes.
(RQ2.2) How does the size of SOLMs impact the performance-resource trade-offs in automated logging? Although we focus on SOLMs, there is still considerable variation in size within this class. Larger models might capture more complex code patterns and lead to higher accuracy, but they could also incur greater computational costs during fine-tuning and inference. Therefore, evaluating the impact of model size is essential to understand the performance-resource trade-offs specific to automated logging.
(RQ2.3) Does the instruct variant of a SOLM outperform its base counterpart for automated logging? The distinction between using a ‘base’ pre-trained model versus an ‘instruct’ version of an SOLM presents another critical strategic choice. Instruct models are tuned to follow user prompts and formats, which is useful for specific tasks, but their tuning is generally broad. Base models, on the other hand, represent the raw capabilities learned during pre-training and might offer greater plasticity when fine-tuned on a highly specific downstream task like ours. Therefore, we aim to clarify which model mode serves as a better foundation for this task.
@@
For RQ2.2, focusing on the influence of model size, we selected the Qwen2.5-coder model series. This choice is driven by the public availability of multiple versions within the same model family, specifically those with 0.5B, 1.5B, 3B, 7B, and 14B parameters, enabling a controlled comparison. Based on the findings from RQ2.1 where LoRA demonstrated superior performance among the PEFT techniques, we exclusively employ LoRA for fine-tuning across these different model sizes. Furthermore, we evaluat each size using both the basic prompt (‘base’) and the RAG-enhanced prompt (‘RAG’), in order to explore whether the effectiveness of RAG varies with model scale, particularly to assess the RAG capabilities of the smaller SOLMs.
To address RQ2.3, we compare the performance of ‘base’ models against their ‘instruct’ counterparts for automated logging. We select the Mistral 7B model, available in both base and instruct variants, for a controlled comparison, as it got the best perfomance in RQ2.1. Both model modes are fine-tuned using LoRA and employed the RAG-enhanced prompt, consistent with the result of RQ2.1 and RQ2.2. The fine-tuning process spanned five epochs, and performance are evaluated using metrics including Reject Rate, PA, LA, AOD, PMR, F1, BLEU-4, and ROUGE-L. The Reject Rate, defined as the percentage of test set examples rejected by the fine-tuned model during inference due to misalignment with learned task-specific criteria, quantifies the model’s selectivity, reflecting its ability to adhere to prompt instructions and generate valid logging statements.
Table 3. Performance Comparison of PEFT Techniques for four 7B SOLMs in automated logging.
@@
Results. (RQ2.1) Table 3 shows the performance comparison of PEFT techniques for SOLMs, where we observe that all evaluated PEFT methods consistently improve performance over the baseline across all SOLMs and most metrics. For instance, looking at the prediction accuracy, QLoRA fine-tuning increased PA from 13.32 to 45.27 for LLaMA3, from 12.85 to 61.90 for Mistral, from 6.99 to 58.97 for CodeLlama, and from 4.30 to 60.21 for Qwen2.5Coder. Similar substantial gains are observed across other metrics like F1 score for variable prediction and BLEU-4/ROUGE-L for statement generation, indicating that fine-tuning with parameter-efficient techniques, is crucial for adapting SOLMs to this specific task.
Furthermore, LoRA emerges as the most effective PEFT methodology, achieving the highest scores for the majority of metrics across all four models. For example, with LoRA, LLaMA3 gets the top PA (56.84), PMR (50.15), F1 (58.22), BLEU-4 (19.84), and ROUGE-L (40.89). Similarly, LoRA leads to the best PA (63.97), LA (69.50), BLEU-4 (23.40), and ROUGE-L (45.73) for Mistral. CodeLlama shows LoRA as the top performer across all metrics. For Qwen2.5Coder, LoRA is also dominant, securing the best PA (62.40), LA (68.46), AOD (88.82), PMR (52.24), BLEU-4 (23.04), and ROUGE-L (44.96). Moreover, QLoRA is competitive to LoRA, achieving the second-best results or even surpassing LoRA in a few specific instances (e.g., F1 for Qwen2.5Coder, PMR, and F1 for Mistral). Prompt Tuning shows some strength, particularly for ‘level’ prediction with LLaMA3, while Prefix Tuning, though an improvement over the baseline, is generally outperformed by LoRA, QLoRA, and Prompt Tuning.
@@
Findings 3: Fine-tuning with PEFT techniques significantly enhances SOLMs performance for automated logging, with LoRA demonstrating the most consistent and superior results across the evaluated models and metrics.
@@
(b) Qwen2.5-coder of Different Sizes
@@
Table 4. Impact of Model Size on Performance and Resource Usage for automated logging.
@@
(RQ2.2) Table 4 reveals that increasing model size leads to improved performance in automated logging, particularly for models 3B and larger. Across almost all metrics, there is a discernible improvement as the model parameter count increases from 0.5B to 14B. For example, prediction accuracy (PA) improves from 21.38 (0.5B) to 66.20 (14B), and ROUGE-L scores for text generation increase from 39.55 (0.5B) to 47.22 (14B). The 14B model consistently outperforms all smaller variants, and the 7B model also shows strong performance.